Thanks very much helping us with the homeworks
1.The sum of 3 consecutive numbers is 93 find the numbers.
2. A father is currently 5 times as old as his son, the difference between their ages is 36 years, how old the son now?
3. Jane has 3782 juice apples for sale in her shop. Fruit and veggies city has 30672 more apples for sale than Jane. How many apples for sale altogether?

Answer by Theo(13342)

n + n + 1 + n + 2 - 93
combine like terms to get:
3n + 3 = 93
subtract 3 from both sides of the equation to get:
3n = 90
solve for n to get:
n = 90/3 = 30
the 3 numbers are 30, 31, 32.
their sum is 93.

f = 5s
f - s = 36
f is the father's age.
s is the son's age.
from the second equation, solve for f to get:
f = s + 36
replace f in the first equation with s + 36 to get:
s + 36 = 5s
subtract s from both sides of the equation to get:
36 = 4s
solve for s to get:
s = 36/4 = 9
the son is 9.
the father is 9 + 36 = 45.
f = 5x becomes 45 = 5*9 which becomes 45 = 45, confirming the value of f and s in this equation is true.
f - s = 36 becomes 45 - 9 = 36, confirming the value of f and s in this equation is also true,
the answer is confirmed to be good.
the solution is that the son is 9 years old now.

jane has 3782 apples for sale.
fruits and veggies has 30672 more apples for sale than jane = 3782 + 30672 = 34454.
together they have 3782 + 34454 = 38236 apples for sale.